2 question (20%) The management of Seligman Manufactoring has decided to tie employee compensation to EVA performance of the firm. The firms CFO, Virginia Whitten, is to make a presentation to the CEO and board of directors illustrating how the program will work under both an unbounded and a bounded plan for awarding incentive compensation for a typical plant manager. To illustrate two plans, Virginia has chosen to use the compensation for a typical plant manager. Under proposed compensation plan, a plant manager would receive a base pay level of $100 000 plus incentive pay equal to 20 percent of this base if the firm hits its EVA performance target exactly Required: Calculate the plant managers incentive pay and total compensation for actual EVA performan 000; 5 20 000 000; and $30 000 000if the target level of EVA performance is set at $20 000 0 Now estimate the plant manag performance and target EVA but with a bounded incentive pay system that has a floor equal to 80 percent of the target performance level and a cap at 120 percent. What are the incentive effects for the plant manager of placing the floor and cap on target performance when determining the incentive pay? What plan(unbounded or bounded) would you recommend to Seligman Corporation and why? a. cedo 15 00n b. ers incentive pay and total compensation for the same three level of EVA
Answer #1

a. For an unbounded incentive pay system

Considering, Target level of EVA performance = $20 000 000 ; Base Pay = $100 000 ; Fraction of pay at risk = 20% or (0.20)

Note: We use the following formula for Unbounded incentive pay system

Incentive pay = (Base pay)*(fraction of pay at risk)*(actual performance/target performance)

Total Compensation = Base Pay + Incentive Pay

Calculation of Incentive pay and Total Compensation:

  • If Actual EVA performance = $15 000 000

Incentive pay = ($100 000)*(0.2)*($15 000 000/$20 000 000) = $15 000

Total Compensation = $100 000 + $15 000 =  $115 000

  • If Actual EVA performance = $20 000 000

Incentive pay = ($100 000)*(0.2)*($20 000 000/$20 000 000) = $20 000

Total Compensation = $100 000 + $20 000 =  $120 000

  • If Actual EVA performance = $30 000 000

Incentive pay = ($100 000)*(0.2)*($30 000 000/$20 000 000) = $30 000

Total Compensation = $100 000 + $30 000 =  $130 000

b. With a bounded incentive pay system that has a floor equal to 80 percent of the target performance level and a cap at 120 percent

Considering, Target level of EVA performance = $20 000 000 ; Base Pay = $100 000 ; Fraction of pay at risk = 20% or (0.20)

Note: Under the 80/120 bounded incentive pay system, incentive is paid only if the actual performance is within 80% and 120% of the target performance. No incentive is paid below 80 percent and above 120 percent of the target performance level.

Floor of Incentive plan i.e. 80 percent of Target performance = (0.80)*($20 000 000) = $16 000 000

Cap of Incentive plan i.e. 120 percent of Target performance = (1.2)*($20 000 000) = $24 000 000

Calculation of Incentive pay and Total Compensation:

  • If Actual EVA performance = $15 000 000

Since, Actual EVA performance (i.e.$15 000 000) is less than the Floor of Incentive plan (i.e.$16 000 000)

Therefore, Incentive pay = $0 and Total Compensation = $100 000

  • If Actual EVA performance = $20 000 000

Since, Actual EVA performance (i.e.$20 000 000) lies within the range

Therefore, Incentive pay = (0.20)*($100 000) = $20 000 and Total Compensation = $100 000 + $20 000 = $120 000

  • If Actual EVA performance = $30 000 000

Since, Actual EVA performance (i.e.$30 000 000) is greater than the Cap of Incentive plan (i.e.$24 000 000)

Therefore, Incentive pay = $0 and Total Compensation = $100 000

The incentive effects for the plant manager of placing the floor and cap on target performance when determining the incentive pay are that it encourages the plant manager to keep his actual EVA performance within the performance range only in order to be eligible for incentive pay.

c. I would recommend the unbounded pay system to Seligman Corporation because in an unbounded system employee incentive is directly linked to the firm performance. The employees will be incentivized to improve firm performance with no floor and cap associated with it. Under the unbounded incentive plan, they won't have to limit themselves to a certain range of performance.